OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 4 No Solution 0 All Real Numbers 1

OpenStudy (anonymous):

All Real Numbers

OpenStudy (anonymous):

4(x+1)=4x+4 4x+4=4x+4 x=x

OpenStudy (anonymous):

thank you, could you help me with like 2 more questions?

OpenStudy (anonymous):

ok, I will do my best

OpenStudy (anonymous):